OpenStudy (anonymous):

An arch of balloons decorates the entrance to a high school prom. The balloons are tied to a frame. The shape of the frame can be modeled by the graph of the equation y=-1/4x^2+3x where x and y are measured in feet. What is the maximum height of the arch of balloons?

OpenStudy (zpupster):

you want to find the max height or vertex ax^2 + bx + c ---quadratic equation x=-b/2a plugging in values gives the vertex x = -(3)/2(-1/4)= ??? now you found x-- plug into the equation to find y(max ht).

OpenStudy (anonymous):

thank you this helps

OpenStudy (anonymous):

I worked problem out like you advised and got X=-1.75, Y=-4.375, did I do this right

OpenStudy (zpupster):

x = -(3)/2(-1/4)= -3/(-2/4)= inversing -3(4)/-2 = 6 y=-1/4(6^2)+3(6)=?? maximum height